OpenStudy (anonymous):

Two cars leave home, driving in opposite directions. One car is 10 miles per hour faster than the other. After 4.5 hours, they are 603 miles apart. How fast are they each driving?

OpenStudy (anonymous):

so one car is going x miles per hour, and the other is going x+10 miles per hour. After 4.5 hours they have traveled 603 miles. so 4.5x + 4.5(x+10) = 603. 4.5(x + x + 10) = 603 2x + 10 = 134 2x = 124 x = 62 (one car) x + 10 = 72 (the other car)

OpenStudy (anonymous):

PS - you can rearrange the letters of your name to spell "big boobs." Is that on purpose?!

OpenStudy (anonymous):

Hahaha, no.